# Safe Navigation under State Uncertainty: Online Adaptation for Robust Control Barrier Functions

### v1

@@ I Introduction
Early robust CBF methods addressed unmodeled dynamics through input-to-state safety margin [4]. Later formulations ensure safety by introducing adaptive robustness margins to reduce conservatism [5, 3]. One recent approach [6] learns a predictive robustness term that is updated online to tighten the CBF inequality whenever the full-order robot dynamics model deviates from its reduced-order model. To enhance robustness in safety constraints, [7] presents a learning-based framework that enables online adaptation of class-$\mathcal{K}_{\infty}$ functions for systems under input constraints.
Beyond model uncertainty, achieving robustness against measurement errors is a notoriously challenging problem because the CBF condition is evaluated on the estimated state. This problem has been approached by using more conservative conditions [8, 9], by considering specific types of state estimators [10, 11], or by adopting stochastic approaches [12]. Although effective in specific control problems, each of these methods imposes specific requirements on the state estimator or uncertainty bounds, which may not be applicable for some real-world applications. To overcome these limitations, a concept for robust CBFs (R-CBFs) is introduced in [13] to ensure safety in the presence of state uncertainty, without requiring prior knowledge of the uncertainty magnitude. Unlike the approach proposed in [11, 10], R-CBFs do not depend on specific classes of state observers. Nevertheless, R-CBFs account for estimation errors through conditions with constant robustness parameters, which may lead to overcompensation for uncertainty, ultimately reducing system performance in an undesirable manner.
@@ II-A Control Barrier Functions
Let $\mathcal{C}(t)\!\subseteq\!\mathcal{X}$} be the 0-superlevel set of a continuously differentiable function $h\!:\!\mathbb{R}_{\geq 0}\!\times\!\mathcal{X}\!\to\!\mathbb{R}$} with 0 as a regular value. The function $h$ is a control barrier function for (1) on $\mathcal{C}(t)$ if there exists an extended class-$\mathcal{K}_{\infty}$ function***A continuous function $\alpha\!:\![0,a)\!\to\!\mathbb{R}_{>0}$}, where a>0{a\!>\!0}, belongs to class-$\mathcal{K}$} ($\alpha\!\in\!\mathcal{K}$}) if it is strictly monotonically increasing and $\alpha(0)\!=\!0$}. A continuous function $\alpha\!:\!\mathbb{R}\!\to\!\mathbb{R}$} belongs to the set of extended class-$\mathcal{K}_{\infty}$ functions ($\alpha\!\in\!\mathcal{K}_{\infty}^{e}$}) if it is strictly monotonically increasing, $\alpha(0)\!=\!0$}, $\lim_{r\!\to\!\infty}\alpha(r)\!=\!\infty$} and $\lim_{r\!\to\!-\infty}\alpha(r)\!=\!-\infty$}. A continuous function $\beta\!:\![0,a)\!\times\!\mathbb{R}^{+}_{0}\!\to\!\mathbb{R}^{+}_{0}$} belongs to class-$\mathcal{KL}$} ($\beta\!\in\!\mathcal{KL}$}), if for every $s\!\in\!\mathbb{R}^{+}_{0}$}, $\beta(\cdot,s)$} is a class-$\mathcal{K}$ function and for every $r\!\in\![0,a)$}, $\beta(r,\cdot)$} is decreasing and $\lim_{s\!\to\!\infty}\beta(r,s)\!=\!0.$} $\alpha\!\in\!\mathcal{K}_{\infty}^{e}$} such that $\forall\mathbf{x}\!\in\!\mathcal{C}(t)$}:
@@ III-A Robust Control Barrier Functions
In practice, feedback control systems rely on an estimate, $\hat{\mathbf{x}}\!\in\!\hat{\mathcal{X}}\!\subseteq\!\mathbb{R}^{n}$} of the true state $x\!\in\!\mathcal{X}\!\subseteq\!\mathbb{R}^{n}$} for the calculation of the control input. This estimate is typically obtained through an observer or filter, and is subject to bounded estimation uncertainties satisfying $\|{\mathbf{x}}\!-\!\hat{\mathbf{x}}\|\!\leq\!\delta({\mathbf{x}})$}, for some $\delta\!:\!\mathcal{X}\!\to\!\mathbb{R}_{\geq 0}$}, where ${\mathbf{x}}\!-\!\hat{\mathbf{x}}\!\triangleq\!\mathbf{e}({\mathbf{x}})$} is the state estimation error function, $\mathbf{e}:\mathcal{X}\!\to\!\mathbb{R}^{n}$}.
@@
Every non-positive super-level set of $h$ is compact, i.e., $\mathcal{C}_{\beta}(t)\!\triangleq\!\big{\{}{\mathbf{x}}\!\in\!{\mathcal{X}}\!:\!h(t,{\mathbf{x}})\!\geq\!-\beta\big{\}}$} is compact for all $\beta\!\in\!\mathbb{R}_{\geq 0},\,t\geq\!0$.
@@
for all $^{\hat{\mathbf{x}}\!\in\!\hat{\mathcal{X}}}$, where $\hat{\mathbf{x}}\!=\!{\mathbf{x}}\!-\!\mathbf{e}({\mathbf{x}})$}.
Given an R-CBF h{h} and a corresponding $\alpha$}, the pointwise set of all control values that satisfy (LABEL:eq:rob_cbf) is given by
@@
We can establish formal robust safety guarantees based on Definition 3 with the help of the following theorem [13]:
@@
The proof follows from Theorem 2 and Corollary 3 in [13], except for the explicit time dependence of $h$, and state-dependent parameters $\gamma_{1},\gamma_{2}$}. ∎
@@
Definition 3 and Theorem 1 extend the R-CBF framework from [13] in two key ways: it employs state-dependent parameters ${\gamma_{1}(\hat{\mathbf{x}}),\gamma_{2}(\hat{\mathbf{x}})}$ rather than constant parameters $\gamma_{1},\gamma_{2}\!\in\!\mathbb{R}_{>0}$}, and it characterizes robust safety for time-varying CBFs. Thus, this definition generalizes Tunable R-CBFs, and Corollary 2 in [13] becomes a special case when the state-dependent parameters take the form $\gamma_{1}(\hat{\mathbf{x}})\!=\!\frac{\gamma_{1}}{\varepsilon_{1}(h(t,\hat{\mathbf{x}}))}$} and $\gamma_{2}^{2}(\hat{\mathbf{x}})\!=\!\frac{\gamma_{2}^{2}}{\varepsilon_{2}(h(t,\hat{\mathbf{x}}))}$}, where $\varepsilon_{1},\varepsilon_{2}\!:\!\mathbb{R}\!\to\!\mathbb{R}_{>0}$} are continuous functions that satisfy $\varepsilon_{i}(r)\!=\!1$} when $r\!\leq\!0$}, and $\frac{d\varepsilon_{i}}{dr}\!>\!0$} when r>0{r\!>\!0}, for i=1,2{i\!=\!1,2}.
@@ III-B Online Parameter Adaptation
$\bar{\gamma}_{1}(t,\!\hat{\mathbf{x}},\!\delta),\bar{\gamma}_{2}(t,\!\hat{\mathbf{x}},\!\delta)\!\!=\!\!\!\operatorname*{arg\,min}_{\gamma_{1},\gamma_{2}\in\mathbb{R}_{>0}}\!\!\!\left(\!\!\frac{\max\{\hat{\sigma}(t,\!\hat{\mathbf{x}},\!\delta,\!\gamma_{1},\!\gamma_{2})\!-\!\gamma_{1},0\}}{2\gamma_{2}}\!\!\right)\!\!.$ (10)
Since it is complex to explicitly compute $\frac{d\hat{\sigma}}{\gamma_{1}},\frac{d\hat{\sigma}}{\gamma_{2}}$, we resort to using derivative-free optimization methods.
@@
To obtain a robust safe control $\mathbf{k}(t,{\mathbf{x}})$, we may replace the nominal CBF condition in the CBF-QP with an R-CBF condition to obtain an R-CBF-QP. When optimizing the parameters $\gamma_{1}$ and $\gamma_{2}$ via (10), this R-CBF-QP needs to be solved for each sampling point $\tilde{\mathbf{x}}_{i}$. When $\mathcal{U}\!=\!\mathbb{R}^{m}$}, closed-form solutions for a QP with single constraint can facilitate fast computation.
If we utilize multiple CBFs, resulting from multiple independent safety specification, directly in the same QP, then we need to optimize two different robustness parameters for each CBF, which consequently increases the dimensionality of the parameter optimization problem. In contrast, by using a single unified CBF constraint in the R-CBF-QP, we only need to optimize a single set of parameters $\bar{\gamma}_{1},\bar{\gamma}_{2}$}, thereby simplifying the optimization process. This motivates the need for a general CBF unification method, as will be described in the next section.
@@ IV-A Safety Filter Construction
Figure 3: Nominal (no state uncertainty): COCP, vanilla CBF-QP, and our R-CBF-QP with online adaptation. Top: Robot trajectories with the safe set and the reference. The brown dot ($\bullet$) marks a vanilla CBF-QP deadlock: the robot reaches a point where no admissible input both preserves safety and moves toward the reference. The baseline controller $\mathbf{k_{d}}$ follows the sinusoidal reference with negligible steady-state error. Bottom: Time histories of $v$, $\omega$, $\gamma_{1},\gamma_{2}$}, and $h$. Our method maintains $h\!\geq\!0$} and closely tracks the reference, similar to COCP. The tuned parameters are obtained as $\bar{\gamma}_{1}(\hat{\mathbf{x}})\!\approx\!0$}, $\bar{\gamma}_{2}(\hat{\mathbf{x}})\!\approx\!0$} with the proposed online tuning method, as $\mathbf{e}({\mathbf{x}})\!\equiv\!0$}.
Figure 4: Robust (bounded state uncertainty): R-COCP, R-CBF-QP with fixed $(\gamma_{1}\!=\!1.4,\gamma_{2}\!=\!0.3)$}, and tunable R-CBF-QP with $\gamma_{1}(h),\gamma_{2}(h)$} given (19), and our R-CBF-QP with online adaptation. Top: Trajectories under the measurement uncertainty. Shaded tubes show the uncertainty envelope from bounded measurement error. Our R-CBF-QP and tunable R-CBF-QP track the reference more closely and are comparable to R-COCP near obstacles, while R-CBF-QP with fixed parameters is more conservative. Bottom: Time histories of $v$, $\omega$, $\gamma_{1},\gamma_{2}$}, and $h$. All methods satisfy $h\!\geq\!0$}. Our online tuning method increases $\bar{\gamma}_{1},\bar{\gamma}_{2}$} only when needed, such as near the boundary.
@@
where $\mu\!\in\!\mathbb{R}_{>0}$} is a tuning parameter and $V\!:\!\mathbb{R}_{\geq 0}\!\times\!\mathcal{X}\!\to\!\mathbb{R}_{\geq 0}$} is a Lyapunov function that characterizes the convergence of $\theta$ to $\theta_{s}(t,x,y)$ under control (13). The function $\theta_{s}\!:\!\mathbb{R}_{\geq 0}\!\times\!\mathbb{R}^{2}\!\to\!\mathbb{S}^{1}$} defines a heading which, when achieved, guarantees safety for the unicycle system. Systematically, $\theta_{s}(t,x,y)$ can be chosen as the direction of the safe output vector for an equivalent integrator system. This modified barrier function (16) has been shown to be effective in eliminating deadlocks when avoiding obstacles in the configuration space [22, Example 6] and [20]. Further details on the construction of $\theta_{s}$ are provided in Appendix -B.
@@
which aims to maintain the desired safe linear velocity while aligning the heading angle towards $\theta_{s}$. Practically, this modification represents the potentially unsafe reference trajectory, $x_{d}(t)~y_{d}(t)]^{\top}$}, being projected onto the safe set.
@@ IV-B Simulations
In the nominal case, the benchmarks are CBF-QP and COCP. For the robust case, we compare the proposed R-CBF-QP with online robustness parameter adaptation ${\bar{\gamma}_{1}(\hat{\mathbf{x}}),\bar{\gamma}_{2}(\hat{\mathbf{x}})}$, R-CBF-QP with fixed $\gamma_{1},\gamma_{2}$} parameters, R-CBF-QP with tunable (CBF-dependent) $\gamma_{1}(h),\gamma_{2}(h)$} parameters, and R-COCP. We use two performance metrics $J_{opt}$ and $J_{t}$ (see Appendix -A) to compare alternative methods. The code and additional details of our simulations can be found at https://github.com/ersindas/robust-CBFs
@@
Next, we choose a sinusoidal reference trajectory tracking task via (15) and (14) such that it yields a safety constraint violation under the nominal controller in $\mathbf{k_{d}}$ in (13). We set the parameters as $v_{\text{ref}}\!=\!0.25$} m/s, $a_{d}\!=\!1.5$} m, $c_{d}\!=\!1.52$} rad/s, and $\varphi_{d}\!=\!0$}. We use $\mathbf{k_{d}}$ and its safe projection $\mathbf{\bar{k}_{d}}$ in (18) with the controller gains $K_{v}\!=\!1,K_{\omega}\!=\!2.5$}. The initial condition is $\mathbf{x}_{0}\!=\![0~-0.35~0]^{\top}$}.
For CBF-QP and R-CBF-QPs, the safety-critical control parameters are: $\alpha\!=\!3$}, and $\mu\!=\!3.3$}. We pick the fixed robustness parameters for comparison as $\gamma_{1}\!=\!1.4$}, $\gamma_{2}\!=\!0.3$}. To modulate the fixed robustness gains in the tunable R-CBF-QP, we utilize an exponential function: $\varepsilon_{i}(h(t,\hat{\mathbf{x}}))={\rm e}^{\left(\eta_{i}\max\{h(t,\hat{\mathbf{x}}),~0\}\right)}$, where $\eta_{i}\!>\!0$}, and in our simulations, we set $\eta_{1}\!=\!2,\eta_{2}\!=\!2$}; therefore, we have
$\gamma_{1}(h)\!=\!\!\frac{\gamma_{1}}{{\rm e}^{\left(\eta_{1}\max\{h(t,\hat{\mathbf{x}}),\!~0\}\right)}},\gamma_{2}^{2}(h)\!=\!\frac{\gamma_{2}^{2}}{{\rm e}^{\left(\eta_{2}\max\{h(t,\hat{\mathbf{x}}),\!~0\}\right)}}.$ (19)
To solve the online tuning problem (10), we sample N=100{N\!=\!100} perturbed states, and perform a gradient-free grid search over $\bar{\gamma}_{1},\bar{\gamma}_{2}$} pairs arranged on a $400\!\times\!400$} mesh covering the range [0.0001,4]{[0.0001,4]} for each axis.
@@
TABLE I: Performance metrics for simulation.
@@
TABLE II: Performance metric $J_{t}$ for hardware.

### v2

@@ I Introduction
Early robust CBF methods addressed unmodeled dynamics through input‑to‑state safety margin [4]. Later formulations ensure safety by introducing adaptive robustness margins to reduce conservatism [5, 3]. One recent approach [6] learns a predictive robustness term that is updated online to tighten the CBF inequality whenever the full-order robot dynamics model deviates from its reduced-order model, assuming accurate state or output measurements. To enhance robustness in safety constraints, [7] presents a learning-based framework that enables online adaptation of class-$\mathcal{K}_{\infty}$ functions for systems under input constraints.
Beyond model uncertainty, robustness to measurement errors is notoriously challenging because CBF conditions are evaluated on the estimated state. This problem has been approached by using more conservative conditions [8, 9], by considering specific types of state estimators [10, 11], or by adopting stochastic approaches [12]. Although effective in specific control problems, each of these methods imposes specific requirements on the state estimator or uncertainty bounds, which may not be applicable to some real-world applications. To overcome these limitations, a concept for robust CBFs (R-CBFs) is introduced in [13] to ensure safety in the presence of state uncertainty, without requiring prior knowledge of the uncertainty magnitude. Unlike the approach proposed in [11, 10], R-CBFs do not depend on specific classes of state observers. Nevertheless, R-CBFs account for estimation errors through conditions with constant robustness parameters, which may lead to overcompensation for uncertainty, ultimately reducing system performance in an undesirable manner.
@@ II-A Control Barrier Functions
Let $\mathcal{C}(t)\!\subseteq\!\mathcal{X}$} be the 0-superlevel set of a continuously differentiable function $h\!:\!\mathbb{R}_{\geq 0}\!\times\!\mathcal{X}\!\to\!\mathbb{R}$} with 0 as a regular value. The function $h$ is a control barrier function for (1) on $\mathcal{C}(t)$ if there exists an extended class-$\mathcal{K}_{\infty}$ function***A continuous function $\alpha\!:\![0,a)\!\to\!\mathbb{R}_{\geq 0}$}, where a>0{a\!>\!0}, belongs to class-$\mathcal{K}$} ($\alpha\!\in\!\mathcal{K}$}) if it is strictly monotonically increasing and $\alpha(0)\!=\!0$}. A continuous function $\alpha\!:\!\mathbb{R}\!\to\!\mathbb{R}$} belongs to the set of extended class-$\mathcal{K}_{\infty}$ functions ($\alpha\!\in\!\mathcal{K}_{\infty}^{e}$}) if it is strictly monotonically increasing, $\alpha(0)\!=\!0$}, $\lim_{r\!\to\!\infty}\alpha(r)\!=\!\infty$} and $\lim_{r\!\to\!-\infty}\alpha(r)\!=\!-\infty$}. A continuous function $\beta\!:\![0,a)\!\times\!\mathbb{R}_{\geq 0}\!\to\!\mathbb{R}_{\geq 0}$} belongs to class-$\mathcal{KL}$} ($\beta\!\in\!\mathcal{KL}$}), if for every $s\!\in\!\mathbb{R}_{\geq 0}$}, $\beta(\cdot,s)$} is a class-$\mathcal{K}$ function and for every $r\!\in\![0,a)$}, $\beta(r,\cdot)$} is decreasing and $\lim_{s\!\to\!\infty}\beta(r,s)\!=\!0.$} $\alpha\!\in\!\mathcal{K}_{\infty}^{e}$} such that $\forall\mathbf{x}\!\in\!\mathcal{C}(t)$}:
@@ III-A Robust Control Barrier Functions
In practice, feedback control systems rely on an estimate, $\hat{\mathbf{x}}\!\in\!\hat{\mathcal{X}}\!\subseteq\!\mathbb{R}^{n}$} of the true state $x\!\in\!\mathcal{X}\!\subseteq\!\mathbb{R}^{n}$} for the calculation of the control input. This estimate is typically obtained through an observer or filter, and is subject to bounded estimation uncertainties satisfying $\|{\mathbf{x}}\!-\!\hat{\mathbf{x}}\|\!\leq\!\delta({\mathbf{x}})$}, for some $\delta\!:\!\mathcal{X}\!\to\!\mathbb{R}_{\geq 0}$}, where $^{\mathbf{e}\!\triangleq\!{\mathbf{x}}\!-\!\hat{\mathbf{x}}}$ is the state estimation error.
@@
Every non-positive super-level set of $h$ is compact and uniformly bounded, i.e., $\mathcal{C}_{\beta}(t)\!\triangleq\!\big\{{\mathbf{x}}\!\in\!{\mathcal{X}}\!:\!h(t,{\mathbf{x}})\!\geq\!-\beta\big\}$ is compact for all $\beta\!\in\!\mathbb{R}_{\geq 0},~t\!\geq\!0$, and $\mathcal{C}_{\beta}(t)\!\subseteq\!S$ for all $t\!\geq\!0$, where $S$ is some compact set.
@@
for all $^{\hat{\mathbf{x}}\!\in\!\hat{\mathcal{X}}}$, where $\hat{\mathbf{x}}\!=\!{\mathbf{x}}\!-\!\mathbf{e}$}.
Compared to the robust CBF formulation in [13], which introduces a general robustness function, our R-CBF definition uses explicit state-dependent robustness terms $\gamma_{1}$ and $\gamma_{2}$. This structured choice is a specialization of the framework in [13] and is tailored to simplify the online adaptation procedure. Given an R-CBF h{h} and a corresponding $\alpha$}, the pointwise set of all control values that satisfy (LABEL:eq:rob_cbf) is given by
@@
We can establish formal robust safety guarantees based on Definition 3 with the help of the following theorem:
@@
The proof follows from Theorem 2 and Corollary 3 in [13], except for the explicit time dependence of $h$, and state-dependent parameters $\gamma_{1},\gamma_{2}$}. The explicit time dependence of $h$ is handled by our assumption that the superlevel sets of $h$ are uniformly bounded in time. ∎
@@
Definition 3 and Theorem 1 extend the R-CBF framework from [13] in two key ways: it employs state-dependent parameters ${\gamma_{1}(\hat{\mathbf{x}}),\gamma_{2}(\hat{\mathbf{x}})}$ rather than constant parameters $\gamma_{1},\gamma_{2}\!\in\!\mathbb{R}_{>0}$}, and it characterizes robust safety for time-varying CBFs. Thus, this definition generalizes tunable R-CBFs [4], and Corollary 2 in [13] becomes a special case when the state-dependent parameters take the form $\gamma_{1}(\hat{\mathbf{x}})\!=\!\frac{\gamma_{1}}{\varepsilon_{1}(h(t,\hat{\mathbf{x}}))}$} and $\gamma_{2}^{2}(\hat{\mathbf{x}})\!=\!\frac{\gamma_{2}^{2}}{\varepsilon_{2}(h(t,\hat{\mathbf{x}}))}$}, where $\varepsilon_{1},\varepsilon_{2}\!:\!\mathbb{R}\!\to\!\mathbb{R}_{>0}$} are continuous functions that satisfy $\varepsilon_{i}(r)\!=\!1$} when $r\!\leq\!0$}, and $\frac{d\varepsilon_{i}}{dr}\!>\!0$} when r>0{r\!>\!0}, for i=1,2{i\!=\!1,2}.
@@ III-B Online Parameter Adaptation
$\bar{\gamma}_{1}(t,\hat{\mathbf{x}},\delta),\bar{\gamma}_{2}(t,\hat{\mathbf{x}},\delta)$
$=\operatorname*{arg\,min}_{\gamma_{1},\gamma_{2}\in\mathbb{R}_{>0}}\left(\frac{\max\{\hat{\sigma}(t,\hat{\mathbf{x}},\delta,\gamma_{1},\gamma_{2})-\gamma_{1},0\}}{2\gamma_{2}}\right).$ (10)
Since it is complex to explicitly compute $\frac{d\hat{\sigma}}{\gamma_{1}},\frac{d\hat{\sigma}}{\gamma_{2}}$, we resort to using derivative-free optimization methods. We remark that $\hat{\sigma}$ in (9) is a practical, sampling-based approximation of $\tilde{\sigma}$.
@@
To obtain a robust safe control $\mathbf{k}(t,{\mathbf{x}})$, we may replace the nominal CBF condition in the CBF-QP with an R-CBF condition to obtain an R-CBF-QP. When optimizing the parameters $\gamma_{1}$ and $\gamma_{2}$ via (10), this R-CBF-QP needs to be solved for each sampling point $\tilde{\mathbf{x}}_{i}$. When $\mathcal{U}\!=\!\mathbb{R}^{m}$}, closed-form solutions for a QP with a single constraint can facilitate fast computation.
If we utilize multiple CBFs, resulting from multiple independent safety specifications, directly in the same QP, then we need to optimize two different robustness parameters for each CBF, which consequently increases the dimensionality of the parameter optimization problem. In contrast, by using a single unified CBF constraint in the R-CBF-QP, we only need to optimize a single set of parameters $\bar{\gamma}_{1},\bar{\gamma}_{2}$}, thereby simplifying the optimization process. This motivates the need for a general CBF unification method, as will be described in the next section.
@@ IV-A Safety Filter Construction
where $\mu\!\in\!\mathbb{R}_{>0}$} is a tuning parameter and $V\!:\!\mathbb{R}_{\geq 0}\!\times\!\mathcal{X}\!\to\!\mathbb{R}_{\geq 0}$} is a Lyapunov function that characterizes the convergence of $\theta$ to $\theta_{s}(t,x,y)$ under control (13). The function $\theta_{s}\!:\!\mathbb{R}_{\geq 0}\!\times\!\mathbb{R}^{2}\!\to\!\mathbb{S}^{1}$} defines a heading which, when achieved, guarantees safety for the unicycle system. Systematically, $\theta_{s}(t,x,y)$ can be chosen as the direction of the safe output vector for an equivalent integrator system. This modified barrier function (16) has been shown to be effective in eliminating deadlocks when avoiding obstacles in the configuration space [22, Example 6] and [20]. Further details on the construction of $\theta_{s}$ are provided in Appendix -B. It is also worth noting that the $\lambda$-superlevel set of $h$ will be contained within the $\lambda$-superlevel set of $h_{0}$ for all $t\geq 0$, for any $\lambda\in\mathbb{R}$, thus preserving compactness of superlevel sets.
@@
which aims to maintain the desired safe linear velocity while aligning the heading angle towards $\theta_{s}$. Practically, this modification projects the potentially unsafe reference trajectory $x_{d}(t)~y_{d}(t)]^{\top}$} onto the safe set, so that the R-CBF-QP only needs to apply small corrections near the boundary, thereby mitigating the jittery behavior.
@@ IV-B Simulations
Figure 3: Nominal (no state uncertainty): COCP, vanilla CBF-QP, and our R-CBF-QP with online adaptation. Top: Robot trajectories with the safe set and the reference. The brown dot ($\bullet$) marks a vanilla CBF-QP deadlock: the robot reaches a point where no admissible input both preserves safety and moves toward the reference. The baseline controller $\mathbf{k_{d}}$ follows the sinusoidal reference with negligible steady-state error. Bottom: Time histories of $v$, $\omega$, $\gamma_{1},\gamma_{2}$}, and $h$. Our method maintains $h\!\geq\!0$} and closely tracks the reference, similar to COCP. The tuned parameters are obtained as $\bar{\gamma}_{1}(\hat{\mathbf{x}})\!\approx\!0$}, $\bar{\gamma}_{2}(\hat{\mathbf{x}})\!\approx\!0$} with the proposed online tuning method, as $\mathbf{e}\!\equiv\!0$}.
Figure 4: Robust (bounded state uncertainty): R-COCP, R-CBF-QP with fixed $(\gamma_{1}\!=\!1.4,\gamma_{2}\!=\!0.3)$}, and tunable R-CBF-QP with $\gamma_{1}(h),\gamma_{2}(h)$} given in (19), and our R-CBF-QP with online adaptation. Top: Trajectories under the measurement uncertainty. Shaded tubes show the uncertainty envelope from bounded measurement error. Our R-CBF-QP and tunable R-CBF-QP track the reference more closely and are comparable to R-COCP near obstacles, while R-CBF-QP with fixed parameters is more conservative. Bottom: Time histories of $v$, $\omega$, $\gamma_{1},\gamma_{2}$}, and $h$. All methods satisfy $h\!\geq\!0$}. Our online tuning method increases $\bar{\gamma}_{1},\bar{\gamma}_{2}$} only when needed, such as near the boundary.
In the nominal case, the benchmarks are CBF-QP and COCP. For the robust case, we compare the proposed R-CBF-QP with online robustness parameter adaptation ${\bar{\gamma}_{1}(\hat{\mathbf{x}}),\bar{\gamma}_{2}(\hat{\mathbf{x}})}$, R-CBF-QP with fixed $\gamma_{1},\gamma_{2}$} parameters, R-CBF-QP with tunable (CBF-dependent) $\gamma_{1}(h),\gamma_{2}(h)$} parameters, and R-COCP. We use two performance metrics $J_{opt}$ and $J_{t}$ (see Appendix -A) to compare alternative methods. The code and additional details of our simulations can be found at
https://github.com/ersindas/robust-CBFs
@@
Next, we choose a sinusoidal reference trajectory tracking task via (15) and (14) such that it yields a safety constraint violation under the nominal controller $\mathbf{k_{d}}$ in (13). We set the parameters as $v_{\text{ref}}\!=\!0.25$} m/s, $a_{d}\!=\!1.5$} m, $c_{d}\!=\!1.52$} rad/s, and $\varphi_{d}\!=\!0$}. We use $\mathbf{k_{d}}$ and its safe projection $\mathbf{\bar{k}_{d}}$ in (18) with the controller gains $K_{v}\!=\!1,K_{\omega}\!=\!2.5$}. The initial condition is $\mathbf{x}_{0}\!=\![0~-0.35~0]^{\top}$}.
For CBF-QP and R-CBF-QPs, the safety-critical control parameters are: $\alpha\!=\!3$}, and $\mu\!=\!3.3$}. We pick the fixed robustness parameters for comparison as $\gamma_{1}\!=\!1.4$}, $\gamma_{2}\!=\!0.3$}. To modulate the fixed robustness gains in the tunable R-CBF-QP, we utilize an exponential function following[4]: $\varepsilon_{i}(h(t,\hat{\mathbf{x}}))={\rm e}^{\left(\eta_{i}\max\{h(t,\hat{\mathbf{x}}),~0\}\right)}$, where $\eta_{i}\!>\!0$}, and in our simulations, we set $\eta_{1}\!=\!2,\eta_{2}\!=\!2$}; therefore, we have
$\begin{split}\gamma_{1}(h)&=\frac{\gamma_{1}}{{\rm e}^{\left(\eta_{1}\max\{h(t,\hat{\mathbf{x}}),~0\}\right)}},\\ \gamma_{2}^{2}(h)&=\frac{\gamma_{2}^{2}}{{\rm e}^{\left(\eta_{2}\max\{h(t,\hat{\mathbf{x}}),\!~0\}\right)}}.\end{split}$ (19)
In this study, we employ a grid search method to solve the online tuning problem (10), as proof-of-concept, though other gradient-free optimizers may also be utilized. We sample N=100{N\!=\!100} perturbed states, and perform a gradient-free grid search over $\bar{\gamma}_{1},\bar{\gamma}_{2}$} pairs arranged on a $400\!\times\!400$} mesh covering the range [0.0001,4]{[0.0001,4]} for each axis.
@@
TABLE I: Performance metrics for simulation in the nominal (no state uncertainty) and robust (bounded uncertainty) scenarios.
@@
Figure 5: Experimental comparison on tracked robot. Pose is measured at the robot’s geometric center; therefore, each circular obstacle boundary is inflated by half the robot’s length. Dashed red circles denote the true safe-set boundaries, while solid red circles indicate the inflated boundaries used in the controller synthesis. Gray tubes visualize the 95% confidence envelope of the $(x,y)$ covariance along the executed paths. (Left) (Top) R-CBF-QP with online adaptation of parameters $\bar{\gamma}_{1},\bar{\gamma}_{2}$}, the proposed method. (Left) (Middle) R-CBF-QP with tunable robustness parameters, i.e., CBF-dependent $\gamma_{1}(h),\gamma_{2}(h)$}. Tunable R-CBF-QP violated the safety constraint in the presence of time-varying state uncertainty, while its performance was similar to R-CBF-QP with online adaptation in simulation. (Left) (Bottom) Robustness parameters for the adaptive and tunable R-CBF-QPs. (Right) (Middle) R-CBF-QP without robustness parameters, i.e., $\gamma_{1}\!=\!0,\gamma_{2}\!=\!0$}. This non-robust controller violates constraints, and the robot falls from the platform and tips over at  24 sec. The experiment was terminated after this unsafe behavior. This result illustrates the necessity of robustification in practice. The CBF-QP–controlled robot encounters a deadlock at the boundary of the first obstacle, marked by the purple cross ( $\times$), since the CBF-QP only uses $v$. And, R-CBF-QP with fixed robustness parameters, i.e., $\gamma_{1}\!=\!1.0,\gamma_{2}\!=\!0.2$}, this controller is more conservative. (Right) (Bottom) The value of CBF $h$ versus time.
IV-C Experimental Results and Comparisons
TABLE II: Performance metric $J_{t}$ for hardware demonstrations.